# Bayesian Additive Regression Trees for functional ANOVA model

### v2

@@ Abstract
Bayesian Additive Regression Trees (BART) is a powerful statistical model that leverages the strengths of Bayesian inference and regression trees. It has received significant attention for capturing complex non-linear relationships and interactions among predictors. However, the accuracy of BART often comes at the cost of interpretability. To address this limitation, we propose ANOVA Bayesian Additive Regression Trees (ANOVA-BART), a novel extension of BART based on the functional ANOVA decomposition, which is used to decompose the variability of a function into different interactions, each representing the contribution of a different set of covariates or factors. Our proposed ANOVA-BART enhances interpretability, preserves and extends the theoretical guarantees of BART, and achieves superior predictive performance. Specifically, we establish that the posterior concentration rate of ANOVA-BART is nearly minimax optimal, and further provides the same convergence rates for each interaction that are not available for BART. Moreover, comprehensive experiments confirm that ANOVA-BART surpasses BART in both accuracy and uncertainty quantification, while also demonstrating its effectiveness in component selection. These results suggest that ANOVA-BART offers a compelling alternative to BART by balancing predictive accuracy, interpretability, and theoretical consistency. The official implementation of ANOVA-BART is publicly available at https://github.com/ParkSeokhun/ANOVA-BART.
@@ 1 Introduction
Bayesian trees and their ensembles have demonstrated significant success in statistics and machine learning (BayesTree ; Bayescart ; BART ; he2019xbart ; lakshminarayanan2015particle ; luo2021bast ; luo2022bamdt ). In particular, Bayesian Additive Regression Trees (BART, BART ), which put the prior mass on the space of ensembles of decision trees and obtain the posterior distribution, have received much attention for their superior prediction performance in various problems including causal inference (hill ; Hahn ), variable selection (Lin ), survival analysis (spara ), interaction detection (due ), smooth function estimation (linero2018bayesian ), mean-variance function estimation (PM ), time series (taddy ), monotone function estimation (mbart ), to name just a few. The theoretical properties of BART (onBART ; bforest ; artbart ; linero2018bayesian ) has been actively explored.
@@
A limitation of BART, however, is that it is a black-box approach in the sense that the relation between inputs and output is hard to be explained. This is because a linear combination of decision trees is not easily interpretable. Recently, interpretability is an important issue in statistics, machine learning and Artificial Intelligence (AI), and constructing interpretable models without hampering prediction performance becomes a critical mission.
@@
Various methods to improve interpretability can be roughly categorized into two approaches - (1) post-processing approach and (2) interpretable model approach. The post-processing approach tries to interpret given black-box models. Representative examples are partial dependency plots (friedman2001greedy ), LIME (ribeiro2016should ) and SHAP (lundberg2017unified ). In contrast, interpretable model approach uses easily interpretable prediction models such as the linear model, generalized additive model (gam ) and more generally functional ANOVA model (func_first ). In particular, the functional ANOVA model has a long history in statistics and received much attention recently in machine learning and AI (func_first ; func_cosso ; func_diag ; func_gaussian ; func_hazard ; func_proj ; MARS ; SSANOVA ; func_hooker ; func_puri ; func_hyper ; func_vae ; park2025tensor ).
The aim of this paper is to develop a Bayesian additive regression trees for the functional ANOVA model, which we call ANOVA-BART. The core idea of ANOVA-BART, which is an interpretable model approach, is to approximate each interaction of the given functional ANOVA model by a linear combination of decision trees. For interpretable modeling with the functional ANOVA model, identifiability must be ensured because explanations are derived from each interaction. Since ordinary decision trees do not satisfy the identifiability condition, we devise special decision trees to ensure the identifiability of each interaction, and propose a prior which is computationally feasible and guarantees an (nearly) minimax optimal posterior concentration rate.
Approximations with ANOVA-BART and BART look similar in the sense that they approximate the true regression function by a linear combination of decision trees. However, the mechanisms of approximation of the two methods are quite different. For BART, the sizes of individual decision trees grow as the sample size increases while the number of trees used in the linear combination is fixed. On the other hand, in ANOVA-BART, the sizes of individual decision trees are fixed that are proportional to the interaction orders while the number of trees grows as the sample size increases. That is, BART approximates the true regression function by a linear combination of few large decision trees while ANOVA-BART does by a linear combination of large numbers of small trees. This qualitative difference makes the MCMC algorithm and technique to derive the posterior concentration rate of ANOVA-BART be much different from those of BART.
@@
This paper is organized as follows. Section 2 reviews the functional ANOVA model and BART. The ANOVA-BART model is described in Section 3, and an MCMC algorithm for the posterior sampling of ANOVA-BART is developed in Section 4. Section 5 provides the posterior concentration rate of ANOVA-BART which is nearly minimax optimal. Section 6 demonstrates that ANOVA-BART outperforms other baseline models by analyzing simulation data as well as multiple benchmark datasets.
@@ 2 Preliminaries
$Y=f(\mathbf{x})+\epsilon,$
where $\mathbf{x}=(x_{1},...,x_{p})^{\top}\in\mathcal{X}\subset\mathbb{R}^{p}$ and $Y\in\mathbb{R}$ are covariate vector and response variable, respectively, $f:\mathcal{X}\rightarrow\mathbb{R}$ is a regression function and $\epsilon\sim N(0,\sigma^{2})$ is a noise. In this section, we review the functional ANOVA model and BART since our proposed method combines these two ideas.
@@ 2.1 Notation
For a real-valued function $f:\mathcal{X}\xrightarrow{}\mathbb{R}$ and $1\leq p<\infty$, we denote $\|f\|_{p,n}:=(\sum_{i=1}^{n}f(\textbf{x}_{i})^{p}/n)^{1/p},$ where $\mathbf{x}_{i}=(x_{i,1},...,x_{i,p})^{\top},i=1,...,n$, are given covariate vectors. We write $\|f\|_{p,\mu}:=(\int_{\mathbf{x}\in\mathcal{X}}f(\mathbb{x})^{p}\mu(d\mathbf{x}))^{1/p},$ where $\mu$ is a probability measure defined on $\mathcal{X}.$ We define $p]=\{1,...,p\}$ to represent the set of integers from $1$ to $p$. For two sequences $a_{n}$ and $b_{n}$, we write $a_{n}=O(b_{n})$ if there exist constants $c>0$ and $n_{0}\in\mathbb{N}$ such that $|a_{n}|\leq c|b_{n}|$ for all $n\geq n_{0}$. For $A=\{a_{1},\ldots,a_{m}\}$ with natural numbers $a_{1}<...<a_{m}$, we write $(x_{j},j\in A):=(x_{a_{1}},...,x_{a_{m}})^{\top}$.
@@ 2.2 Functional ANOVA model
We assume that $\mathcal{X}=\prod_{j=1}^{p}\mathcal{X}_{j}$, where $\mathcal{X}_{j}$ is a subset of $\mathbb{R}.$ For $\mathbb{x}\in\mathcal{X}$ and $S\subseteq[p]$, we write $\mathbb{x}_{S}:=(x_{j},j\in S)$ and $\mathcal{X}_{S}:=\prod_{j\in S}\mathcal{X}_{j}$. For a given function $f_{S}$ defined on $\mathcal{X}_{S}$ and a probability measure $\mu$ on $\mathcal{X}$, we say that $f_{S}$ satisfies the $\mu$-identifiability condition (func_hooker ; func_diag ) if the following holds:
@@
Theorem 2.1 (func_diag ; mcbook ).
@@
The functional ANOVA model with only the main effects is the generalized additive model (GAM, gam ) which is widely used in practice. For machine learning applications, func_hooker ; func_puri ; park2025tensor emphasize the usefulness of the functional ANOVA model as an interpretable AI tool. There are various algorithms to estimate the interactions in the functional ANOVA model including MARS (MARS ), SSANOVA (SSANOVA ), COSSO (func_cosso ), NAM (nam ), ANOVA-TPNN (park2025tensor ) and so on.
@@ 2.3 Bayesian Additive Regression Trees (BART)
BART (BART ) models the regression function $f$ by the sum of random decision trees. That is, BART assumes a priori that $f(\cdot)=\beta_{0}+\sum_{t=1}^{T}\mathbb{T}_{t}(\cdot),$ where $\mathbb{T}_{t}$s are random decision trees and obtains the posterior distribution of $\mathbb{T}_{t}$s. Each decision tree is represented by the three parameters at each node - (i) the binary indicator whether a node is internal or terminal, (ii) the pair of (split-variable, split-criterion) if the node is internal and (iii) the height (i.e the prediction value) if the node is terminal. BART puts a prior mass on these three parameters to generate independent random decision trees $\mathbb{T}_{t}$s while $T$ is fixed in advance and not inferred. Then, BART samples from the posterior distribution $p(\mathbb{T}_{1},\ldots,\mathbb{T}_{T},\sigma^{2}|{\rm data})$ using the backfitting MCMC algorithm (BART ).
@@ 3 ANOVA-BART : BART on the functional ANOVA model
In this section, we propose a version of BART for the functional ANOVA model which we call ANOVA-BART. The main idea of ANOVA-BART is to approximate each component of the functional ANOVA model (2) using a sum of decision trees that satisfy the identifiability condition. That is, we set
$f_{S}(\mathbf{x}_{S})=\sum_{t=1}^{T_{S}}\mathbb{T}_{t}^{S}(\mathbf{x}_{S}),$ (3)
where $\mathbb{T}_{t}^{S}$s are specially designed decision trees defined on $\mathcal{X}_{S}$. Let $\mathcal{T}^{S}$ be the class of decision trees where $\mathbb{T}_{t}^{S}$s belong. In the following subsections, we propose $\mathcal{T}^{S}$ so that $f_{S}(\cdot)$ always satisfies the identifiability condition and devise a prior on $\mathcal{T}^{S}$ for $S\subseteq[p]$ and the other parameters including the numbers of decision trees $T_{S}$ for each $S$ as well as the variance $\sigma^{2}$ of noises. Note that ANOVA-BART treats the number of decision trees as random and also infers it while BART fixes it in advance, which makes ANOVA-BART differs theoretically as well as algorithmically from BART.
@@ 3.1 Choice of $\mathcal{T}^{S}$: Identifiable binary-product trees
A technical difficulty in choosing $\mathcal{T}^{S}$ is that $f_{S}$ should satisfy the identifiability condition (1). To ensure this restriction, we make each decision tree in $\mathcal{T}^{S}$ satisfy the identifiability condition. Let $\mu_{n,S}$ be an empirical distribution of given $\mathbf{x}_{1,S},...,\mathbf{x}_{n,S}$, where $\mathbf{x}_{i,S}=(x_{i,j},j\in S)$ for i=1,,ni=1,...,n. When $S=\{\ell\}$, we write $\mu_{n,\ell}$ instead of $\mu_{n,\{\ell\}}$ for notational simplicity. For the main effects (i.e. $|S|=1$): We first consider decision trees for the main effect. For $\mathbb{T}^{S}\in\mathcal{T}^{S}$ with $S=\{\ell\}$, we consider a tree with one split by the covariate $x_{\ell}.$ That is, when $x_{\ell}$ is continuous covariate, any $\mathbb{T}^{\{\ell\}}\in\mathcal{T}^{\{\ell\}}$ is given as
$\mathbb{T}^{\{\ell\}}(x_{\ell})=\beta_{-1}\mathbb{I}(x_{\ell}\leq s_{\ell})+\beta_{1}\mathbb{I}(x_{\ell}>s_{\ell})$
for $s_{\ell}\in\mathcal{A}_{\ell},$ where $\mathcal{A}_{\ell}$ is the set of all possible split values for $x_{\ell}$ which will be specified later on. To ensure the identifiability condition, we let $\beta_{-1}$ and $\beta_{1}$ satisfy
$\beta_{-1}\mu_{n,\ell}\{x_{\ell}\leq s_{\ell}\}+\beta_{1}\mu_{n,\ell}\{x_{\ell}>s_{\ell}\}=0.$
Note that only one of $\beta_{-1}$ and $\beta_{1}$ is a free parameter and the other is determined automatically by the identifiability condition. Thus, the degree of freedom of $(\beta_{-1},\beta_{1})$ is 1.
For categorical covariate $x_{\ell},$ a decision tree with one split by the covariate $x_{\ell}$ is defined similarly by replacing $\mathbb{I}(x_{\ell}\leq s_{\ell})$ and $\mathbb{I}(x_{\ell}>s_{\ell})$ with $\mathbb{I}(x_{\ell}\in A)$ and $\mathbb{I}(x_{\ell}\in A^{c})$ for a subset $A$ of $\text{supp}(x_{\ell}),$ where $\text{supp}(x_{\ell})$ is the support of $x_{\ell}$. For notational convenience, we only consider continuous covariates in the followings, but all the arguments can be modified easily for categorical covariates.
For the second order interactions (i.e. $|S|=2$): For the second order interaction, we consider the following form of a decision tree for $\mathbb{T}^{\{\ell,k\}}\in\mathcal{T}^{\{\ell,k\}}$:
$\begin{split}\mathbb{T}^{\{\ell,k\}}(x_{\ell},x_{k})=&\beta_{-1,-1}\mathbb{I}(x_{\ell}\leq s_{\ell},x_{k}\leq s_{k})+\beta_{-1,1}\mathbb{I}(x_{\ell}\leq s_{\ell},x_{k}>s_{k})\\ &+\beta_{1,-1}\mathbb{I}(x_{\ell}>s_{\ell},x_{k}\leq s_{k})+\beta_{1,1}\mathbb{I}(x_{\ell}>s_{\ell},x_{k}>s_{k}),\end{split}$ (4)
[…]
$\begin{split}\mu_{n,\ell}\{X_{\ell}\leq s_{\ell}\}\beta_{-1,-1}+\mu_{n,\ell}\{X_{\ell}>s_{\ell}\}\beta_{1,-1}&=0,\\ \mu_{n,\ell}\{X_{\ell}\leq s_{\ell}\}\beta_{-1,1}+\mu_{n,\ell}\{X_{\ell}>s_{\ell}\}\beta_{1,1}&=0,\\ \mu_{n,k}\{X_{k}\leq s_{k}\}\beta_{-1,-1}+\mu_{n,k}\{X_{k}>s_{k}\}\beta_{-1,1}&=0,\\ \mu_{n,k}\{X_{k}\leq s_{k}\}\beta_{1,-1}+\mu_{n,k}\{X_{k}>s_{k}\}\beta_{1,1}&=0.\end{split}$ (5)
Similarly to the main effects, the degree of freedom of ($\beta_{-1,-1},\beta_{-1,1},\beta_{1,-1},\beta_{1,1}$) is 1, that is, one of $\beta_{-1,-1},\beta_{-1,1},\beta_{1,-1},\beta_{1,1}$ determines the other three. See Section A of Supplementary Material for details. For higher order interactions (i.e. $|S|>2$): The decision tree in (4) for $|S|=2$ can be rewritten as
$\mathbb{T}^{\{\ell,k\}}(x_{\ell},x_{k})=\sum_{(v_{\ell},v_{k})\in\{-1,1\}^{2}}\beta_{v_{\ell},v_{k}}\mathbb{I}^{(v_{\ell})}(x_{\ell}-s_{\ell}>0)\mathbb{I}^{(v_{k})}(x_{k}-s_{k}>0)$
with $\mathbb{I}^{(1)}(\cdot):=\mathbb{I}(\cdot)$ and $\mathbb{I}^{(-1)}(\cdot):=1-\mathbb{I}(\cdot)$. We generalize (4) for $S=\{\ell_{1},\ldots,\ell_{d}\}$ with $d>2$ by
$\mathbb{T}^{S}(\mathbf{x}_{S})=\sum_{\mathbf{v}\in\{-1,1\}^{d}}\beta_{\mathbf{v}}\prod_{j=1}^{d}\mathbb{I}^{(v_{\ell_{j}})}(x_{\ell_{j}}-s_{\ell_{j}}>0),$ (6)
where $\mathbf{v}:=(v_{j},j\in S)$. We call a decision tree having the form of (6) a binary-product tree, which has been considered by prokhorenkova2018catboost ; popov2019neural . Binary means that the binary split is used for each covariate and product means that the split rules are produced by the product of the split rules of each covariate. Figure 1 draws the examples of binary-product trees for $|S|=1,2$ and 3.
Similar with (5), the binary-product tree in (6) satisfies the identifiability conditions if and only if the coefficients $\beta_{\mathbf{v}}$ for $\mathbf{v}\in\{-1,1\}^{d}$ satisfy the following conditions: for any $k=1,\ldots,d,$ and any $v_{j}\in\{-1,1\}$ for $j\in S\setminus\{\ell_{k}\}$,
$\sum_{v_{\ell_{k}}\in\{-1,1\}}\beta_{\mathbf{v}}\mu_{n,\ell_{k}}^{(v_{\ell_{k}})}\{X_{\ell_{k}}-s_{\ell_{k}}>0\}=0,$ (7)
where $\mu_{n,\ell_{k}}^{(1)}:=\mu_{n,\ell_{k}}$ and $\mu_{n,\ell_{k}}^{(-1)}:=1-\mu_{n,\ell_{k}}$. That is, Condition (7) extends Condition (5) to the case of general $S$. We refer to such a tree as an identifiable binary-product tree.
@@
The degree of freedom of $(\beta_{\mathbf{v}},\mathbf{v}\in\{-1,1\}^{d})$ is always 1 for every $d\in[p]$ (See Section A of Supplementary Material), and hence the heights of any identifiable binary-product tree are uniquely defined by choosing one of the heights. In this paper, we set $\beta_{-1,\ldots,-1}$ as the free parameter. Then, any identifiable binary-product tree for the component $S$ is parameterized by the vector of the split values $\mathbf{s}:=(s_{j},j\in S)$ corresponding to each covariate in $S$ and the height $\beta_{-1,\ldots,-1}.$ Whenever needed, we write a given identifiable binary-product tree $\mathbb{T}^{S}(\mathbf{x}_{S})$ as $\mathbb{T}(\mathbf{x}:S,\mathbf{s},\beta)$, where $\beta:=\beta_{-1,\ldots,-1}.$

### v3

@@ Abstract
Bayesian Additive Regression Trees (BART) is a powerful statistical model that leverages the strengths of Bayesian inference and regression trees. It has received significant attention for capturing complex non-linear relationships and interactions among predictors. However, the accuracy of BART often comes at the cost of interpretability. To address this limitation, we propose ANOVA Bayesian Additive Regression Trees (ANOVA-BART), a novel extension of BART based on the functional ANOVA decomposition, which is used to decompose the variability of a function into different interactions, each representing the contribution of a different set of covariates or factors. Our proposed ANOVA-BART enhances interpretability, preserves and extends the theoretical guarantees of BART, and achieves comparable prediction performance. Specifically, we establish that the posterior concentration rate of ANOVA-BART is nearly minimax optimal, and further provides the same convergence rates for each interaction that are not available for BART. Moreover, comprehensive experiments confirm that ANOVA-BART is comparable to BART in both accuracy and uncertainty quantification, while also demonstrating its effectiveness in component selection. These results suggest that ANOVA-BART offers a compelling alternative to BART by balancing predictive accuracy, interpretability, and theoretical consistency. The official implementation of ANOVA-BART is publicly available at https://github.com/ParkSeokhun/ANOVA-BART.
@@ 1 Introduction
Bayesian trees and their ensembles have demonstrated significant success in statistics and machine learning (Chipman et al. (1998); Dension et al. (1998); Chipman et al. (2010); He et al. (2019); Lakshminarayanan et al. (2015); Luo et al. (2021, 2022)). In particular, Bayesian Additive Regression Trees (BART, Chipman et al. (2010)), which put the prior mass on the space of ensembles of decision trees and obtain the posterior distribution, have received much attention for their superior prediction performance in various problems including causal inference (Hill (2011); Hahn et al. (2020)), variable selection (Linero (2018a)), survival analysis (Sparapani et al. (2016)), interaction detection (Du and Linero (2019)), smooth function estimation (Linero and Yang (2018)), mean-variance function estimation (Pratola et al. (2020)), time series (Taddy et al. (2011)), monotone function estimation (Chipman et al. (2022)), to name just a few. In addition, theoretical properties of BART have been actively studied (Ročková and Saha (2019); Ročková and van der Pas (2020); Jeong and Ročlková (2023); Linero and Yang (2018))
@@
A limitation of BART, however, is that it is a black-box approach in the sense that the relation between covariates and response variable is hard to be explained. This is because a linear combination of decision trees is not easily interpretable even though each decision tree is interpretable. Recently, interpretability is an important issue in statistics, machine learning and Artificial Intelligence (AI), and constructing interpretable models without hampering prediction performance becomes a key challenge.
@@
Various methods to improve interpretability can be roughly categorized into two approaches - (1) post-processing approach and (2) interpretable model approach. The post-processing approach tries to interpret a given black-box model. Representative examples are partial dependency plots (Friedman (2001)), LIME (Ribeiro et al. (2016)) and SHAP (Lundberg and Lee (2017)). In contrast, the interpretable model approach uses easily interpretable prediction models such as the linear model, generalized additive model (Hastie and Tibshirani (1986)) and more generally functional ANOVA model (Hoeffding (1948)). In particular, the functional ANOVA model has a long history in statistics and received much attention recently in machine learning and AI (Hoeffding (1948); Lin and Zhang (2006); Hooker (2007); Kaufman and Sain (2010); Huang et al. (2000); Huang (1998); Friedman (1991); Gu and Wahba (1993); Hooker (2004); Lengerich et al. (2020); Hutter et al. (2014); Märtens and Yau (2020); Park et al. (2025b)).
The aim of this paper is to develop a Bayesian inference procedure for the functional ANOVA model, which we call ANOVA-BART. The core idea of ANOVA-BART, which is an interpretable model approach, is to approximate each interaction in the functional ANOVA model by a linear combination of decision trees. A technical difficulty is that each component in the functional ANOVA model is not identifiable and thus direct adoption of BART to the functional ANOVA model would result in unstable inference of each component. Similar phenomena are observed in neural networks for the functional ANOVA model such as Neural Additive Models (NAM, Agarwal et al. (2021)) and Neural Basis Models (NBM, Radenovic et al. (2022)).
To resolve this problem, we impose the identifiability condition (See Section 2.2 for details) on each component. Then, we devise special decision trees such that (1) they satisfy the identifiability condition, (2) a linear combination of them approximates a smooth function well to guarantees a (nearly) minimax optimal posterior concentration rate, and (3) an efficient MCMC algorithm can be developed without much difficulty.
Special neural networks to ensure the identifiability of each component in the functional ANOVA model have been developed by Park et al. (2025b) and Park et al. (2025a). As usual, however, neural networks are highly vulnerable to outliers in covariates and thus special preprocessing of covariates are indispensable for stable inference. In contrast, decision trees are robust to outliers and so yield stabler and more accurate prediction models without requiring data preprocessing.
@@
This paper is organized as follows. Section 2 reviews the functional ANOVA model and BART. ANOVA-BART is described in Section 3, and an MCMC algorithm for the posterior sampling of ANOVA-BART is developed in Section 4. Section 5 provides the posterior concentration rate of ANOVA-BART which is nearly minimax optimal and adaptive to the interaction structure (i.e., the highest order of the signal components) as well as the smoothness of the true model. Section 6 demonstrates the competitiveness of ANOVA-BART against baseline models through analyses of simulation data and multiple benchmark datasets.
@@ 2 Preliminaries
$Y|\mathbf{x}\sim\mathbb{P}_{f(\mathbf{x}),\psi}$
where $\mathbf{x}=(x_{1},...,x_{p})^{\top}\in\mathcal{X}\subset\mathbb{R}^{p}$ and $Y\in\mathbb{R}$ are covariate vector and response variable, respectively, $f:\mathcal{X}\rightarrow\mathbb{R}$ is a regression function and $\psi\in\mathbb{R}$ is a nuisance parameter.
We assume that the distribution $\mathbb{P}_{f(\mathbf{x}),\psi}$ belongs to the exponential family and admits a probability density function $p_{f(\mathbf{x}),\psi}$, defined as
$p_{f(\mathbf{x}),\psi}(y)=\exp\bigg({f(\mathbf{x})y-A(f(\mathbf{x}))\over\psi}+S(y,\psi)\bigg),$ (1)
where $A(\cdot)$ is the log-partition function, ensuring that the density integrates to one.
In this section, we review the functional ANOVA model and BART since our proposed method is closely related to these two models.
@@ 2.1 Notation
For a real-valued function $f:\mathcal{X}\xrightarrow{}\mathbb{R}$ and $1\leq p<\infty$, we denote $\|f\|_{p,n}:=(\sum_{i=1}^{n}f(\textbf{x}_{i})^{p}/n)^{1/p},$ where $\mathbf{x}_{i}=(x_{i,1},...,x_{i,p})^{\top},i=1,...,n$, are given covariate vectors. We write $\|f\|_{p,\mu}:=(\int_{\mathbf{x}\in\mathcal{X}}f(\mathbb{x})^{p}\mu(d\mathbf{x}))^{1/p},$ where $\mu$ is a probability measure defined on $\mathcal{X}.$ We define $p]=\{1,...,p\}$ to represent the set of integers from $1$ to $p$. For two sequences $a_{n}$ and $b_{n}$, we write $a_{n}=O(b_{n})$ if there exist constants $c>0$ and $n_{0}\in\mathbb{N}$ such that $|a_{n}|\leq c|b_{n}|$ for all $n\geq n_{0}$. For $A=\{a_{1},\ldots,a_{m}\}$ with natural numbers $a_{1}<...<a_{m}$, we write $(x_{j},j\in A):=(x_{a_{1}},...,x_{a_{m}})^{\top}$. Let $\text{power}([p],d)$ be the collection of all subsets of $[p]$ whose cardinality is $d$.
@@ 2.2 Functional ANOVA model
We assume that $\mathcal{X}=\prod_{j=1}^{p}\mathcal{X}_{j}$, where $\mathcal{X}_{j}$ is a subset of $\mathbb{R}.$ For $\mathbb{x}\in\mathcal{X}$ and $S\subseteq[p]$, we write $\mathbb{x}_{S}:=(x_{j},j\in S)$ and $\mathcal{X}_{S}:=\prod_{j\in S}\mathcal{X}_{j}$. For a given function $f_{S}$ defined on $\mathcal{X}_{S}$ and a probability measure $\mu$ on $\mathcal{X}$, we say that $f_{S}$ satisfies the $\mu$-identifiability condition (Hooker (2004, 2007)) if the following holds:
@@
Theorem 2.1 (Hooker (2007); Owen (2013)).
@@
The functional ANOVA model with only the main effects is the generalized additive model (GAM, Hastie and Tibshirani (1986)) which is widely used in practice. For machine learning applications, Hooker (2004); Lengerich et al. (2020); Park et al. (2025b) emphasize the usefulness of the functional ANOVA model as an interpretable AI tool. There are various algorithms to estimate the interactions in the functional ANOVA model including MARS (Friedman (1991)), SSANOVA (Gu and Wahba (1993)), COSSO (Lin and Zhang (2006)), NAM (Agarwal et al. (2021)), ANOVA-TPNN (Park et al. (2025b)) and so on.
@@ 2.3 Bayesian Additive Regression Trees (BART)
BART (Chipman et al. (2010)) assumes a Gaussian regression model, i.e., $Y|\mathbf{x}\sim N(\cdot|f(\mathbf{x}),\sigma^{2})$, where the regression function $f$ is modeled as the sum of random decision trees. That is, BART assumes a priori that $f(\cdot)=\sum_{t=1}^{T}\mathbb{T}_{t}(\cdot),$ where $\mathbb{T}_{t}$s are random decision trees and obtains the posterior distribution of $\mathbb{T}_{t}$s. Each decision tree is represented by the three parameters at each node - (i) the binary indicator whether a node is internal or terminal, (ii) the pair of (split-variable, split-criterion) if the node is internal and (iii) the height (i.e the prediction value) if the node is terminal. BART puts a prior mass on these three parameters to generate independent random decision trees $\mathbb{T}_{t}$s while $T$ is fixed in advance and not inferred. Then, BART samples from the posterior distribution $p(\mathbb{T}_{1},\ldots,\mathbb{T}_{T},\sigma^{2}|{\rm data})$ using the backfitting MCMC algorithm (Chipman et al. (2010)).
Building upon BART, several extensions have been proposed. Specifically, Linero (2018b) proposed Dirichlet Additive Regression Trees (DART), which equip BART with a Dirichlet prior on the splitting probabilities over covariates, thereby encouraging sparse and adaptive variable selection. Linero and Yang (2018) proposed Soft BART (SBART) by replacing the indicator functions in BART decision trees with sigmoid functions, and proves that the posterior concentration rate of SBART is minimax optimal over Hölder function spaces with smoothness parameter $\alpha\in(0,\infty)$. Deshpande et al. (2024) introduced Varying Coefficient Bayesian Additive Regression Trees (VCBART), which use BART to estimate coefficient functions in varying coefficient models. Recently, Linero (2025) proposed a reversible jump Markov chain Monte Carlo (RJMCMC) algorithm for Generalized BART (GBART) when the data distribution belongs to the exponential family.
@@ 3 ANOVA-BART : BART on the functional ANOVA model
In this section, we propose a version of BART for the functional ANOVA model which we call ANOVA-BART. The main idea of ANOVA-BART is to approximate each component of the functional ANOVA model by the ensemble of random decision trees. That is, we set
$f_{S}(\mathbf{x}_{S})=\sum_{t=1}^{T_{S}}\mathbb{T}_{t}^{S}(\mathbf{x}_{S}),$
where $\mathbb{T}_{t}^{S}$s are specially designed decision trees defined on $\mathcal{X}_{S}$. Let $\mathcal{T}^{S}$ be the class of decision trees where $\mathbb{T}_{t}^{S}$s belong. In the following subsections, we propose $\mathcal{T}^{S}$ so that $f_{S}(\cdot)$ always satisfies the identifiability condition and devise priors on $\mathcal{T}^{S}$ for $S\subseteq[p]$ and the other parameters including the numbers of decision trees $T_{S}$ as well as the nuisance parameter $\psi$.
@@ 3.1 Choice of $\mathcal{T}^{S}$: Identifiable binary-product trees
A technical difficulty in choosing $\mathcal{T}^{S}$ is that $f_{S}$ should satisfy the identifiability condition in (2). To ensure this restriction, we make each decision tree in $\mathcal{T}^{S}$ satisfy the identifiability condition. Let $\mu_{n,S}$ be an empirical distribution of given $\mathbf{x}_{1,S},...,\mathbf{x}_{n,S}$, where $\mathbf{x}_{i,S}=(x_{i,j},j\in S)$ for i=1,,ni=1,...,n. When $S=\{\ell\}$, we write $\mu_{n,\ell}$ instead of $\mu_{n,\{\ell\}}$ for notational simplicity. We let $\mathcal{T}^{S}$ consist of decision trees having the following form
$\mathbb{T}^{S}(\mathbf{x}_{S})=\beta\prod_{j\in S}\big(\mathbb{I}(x_{j}\leq s_{j})+\mathfrak{a}_{j}\mathbb{I}(x_{j}>s_{j})\big),$ (4)
where $\mathfrak{a}_{j}=-\mu_{n,j}\{X_{j}\leq s_{j}\}/\mu_{n,j}\{X_{j}>s_{j}\}$. Note that $\mathbb{T}^{S}$ satisfies the identifiability condition in (2). We refer to such a decision tree as an identifiable binary-product tree.
@@
Figure 1 shows three indentifiable binary-product trees of the orders 1,2 and 3. The decision trees have a special structure: all terminal (leaf) nodes have the same depth and all nodes at the same level share an identical splitting rule. The heights of each terminal node are automatically decided when the height $\beta$ of the left-most terminal node is selected. To sum up, any identifiable binary-product tree is parameterized by the componet set $S,$ the vector of the split values $\mathbf{s}:=(s_{j},j\in S)$ corresponding to each covariate in $S$ and the height $\beta.$ Whenever needed, we write a given identifiable binary-product tree $\mathbb{T}^{S}(\mathbf{x}_{S})$ as $\mathbb{T}(\mathbf{x}:S,\mathbf{s},\beta)$.
For categorical covariate $x_{\ell},$ we proprocess all categorical covariates using the one-hot encoding. That is, each categorical covariate is transformed to a binary vector. Therefore, splits on categorical covariates can be handled similarly to continuous covariates, and we only consider continuous covariates for the remainder of the paper.